In running a 400 meter race, an athlete toes out on each stride so that her foot length measured in the direction of progression is 2.0 centimeters shorter than if she toed straight ahead. If her stride length averaged 2.00 meters per stride, how much more distance could she cover in the same length of time (50 seconds) if she toed straight ahead?
Answer #1

Answer:- 0.8 meter

Explanation:- following data are given in question

Total race = 400 meter

Foot length = 2.0 cm shorter than stride

Average = 2.00 meters per stride

Time = 50 seconds

Distance covered = ???

Solution:-

= 400/50 = 8

= 8/2 = 4

= 4×2.2 = 8.8

= 4×2 = 8

= 0.8 meter Answer